# Seeing Straight: Document Orientation Detection for Efficient OCR

### v1

@@ 5 Experimental Design
In this section, we detail the experimental strategy used to evaluate both the intrinsic accuracy and downstream impact of the proposed rotation correction module. More technical details are provided in the Appendix A.
Baselines: We benchmark against both conventional OCR models and Vision-Language Models (VLMs), corresponding prompts in Table 5. Specifically, we consider docTR liao2023doctr, transformer-based TrOCR li2021trocr and more recent VLMs including open-weights Gemma-3 (27B) deepmind2024gemma, Nanonets-OCR-s nanonets2024ocr, LLaMA-4-17B (Maverick)  touvron2024llama4 supporting Indian languages, alongwith closed-source GPT-4o hurst2024gpt4o, Gemini-2.5 Pro and Flash team2023gemini. Inference for closed-source VLMs was conducted via respective public APIs.
Evaluation metrics: We benchmark against field level accuracy on the SROIE dataset. We adopt the evaluation strategy used in prior and contemporary studies (fu2024ocrbench; anonymous2025multimodal), reporting Average Normalized Levenshtein Similarity (ANLS) as an interpretable metric  (biten2019scene) in addition to other variants like Word Error Rate (WER) and Character Error Rate (CER) (smith2007overview; neudecker2021survey) based on Levenshtein Distance (lcvenshtcin1966binary).
  • Ideal Upright Conditions: OCR on upright images with an ideal canonical setup having no rotation.
Model vs Accuracy ORB-En $\uparrow$ ORB-Indic $\uparrow$
TrOCR 26.22 68.90
docTR 63.28 46.95
Nanonets-OCR-s 24.83 37.50
Gemma-3 27B 30.57 33.23
LLaMA-4 (Maverick) 17B 65.49 56.70
GPT-4o 59.58 50.60
Gemini-2.5 Flash 39.55 41.46
Ours 96.81 92.68
Table 2: Benchmark comparison of our method against 0-shot VLMs and other baselines for document rotation classification (in %) on ORB datasets. Notably, classical docTR and TrOCR perform better than closed-source GPT-4o and Gemini-2.5 Flash on ORB-En (both subsets) and ORB-Indic respectively. Best scores bold, second-best underlined.
Model vs Accuracy
Ideal upright conditions $\uparrow$
w/o Rotation $\uparrow$
w/ Rotation $\uparrow$
@@
Table 3: Field-level OCR accuracy (in %) on ORB-En-SROIE with and w/o our rotation pipeline. We find 2x and 4x improvement in the open-source Tesseract and docTR OCR engines respectively (second best scores underlined).
@@ 6 Results and Discussion
Rotation Classification: Table 2 reports the rotation classification accuracy of all evaluated models on our dataset. Our proposed classifier, comprising approximately 304M parameters, with 0.5s latency on NVIDIA H100, achieves 96.81% accuracy on ORB-En and 92.68% on the multilingual ORB-Indic dataset, demonstrating strong generalization across diverse scripts and layouts. We also conducted an ablation study where dynamic cropping with multi-layer classifier showed the best performance (Table 6 and Section B in Appendix).
Conventional OCR systems such as docTR and TrOCR show inconsistent performance when repurposed for rotation classification. Notably, all VLMs underperformed significantly on this task, reinforcing recent findings that VLMs often struggle with standard computer vision benchmarks (zhang2024visually; cooper2025rethinking; ramachandran2025well). Despite their impressive generative capabilities, models like Gemma-3, LLaMA-4, and GPT-4o failed to reliably infer document orientation. Among closed-source models, GPT-4o slightly outperforms Gemini-2.5 Flash on rotation classification, though both fall well short of our model.
Evaluation on structured OCR: We assess the effect of our rotation correction module on downstream OCR performance using ORB-En-SROIE (Table 3). Field-level accuracy is calculated across four key fields, each weighted equally (25%). Incorporating our module yields substantial improvements: Tesseract doubles its accuracy, and docTR improves fourfold. Interestingly, GPT-4o, despite rotation correction, still underperforms compared to conventional OCR engines on structured field extraction (with 6% gain). Nanonets showed the best performance achieving a 8% gain when used with our module. Gemini-2.5 Pro benefits with a 14% uplift, while Gemini-2.5 Flash shows minimal improvement, suggesting that it may already implement internal rotation handling, either through pre-processing or prior exposure to rotated data during multi-modal training.
@@
ORB-En-SynthDog dataset
docTR 56.23 40.04 88.03 69.07 56.23 40.04
Nanonets-OCR-s 2.61 1.73 44.14 34.25 2.61 1.73
Gemma-3 27B 61.56 30.29 75.15 50.06 61.56 30.29
LLaMA-4 (Maverick) 17B 29.37 14.09 59.91 41.75 29.37 14.09
GPT-4o 82.22 73.65 92.35 82.08 82.22 73.65
Gemini-2.5 Pro 58.58 52.34 70.96 65.41 67.07 61.16
Gemini-2.5 Flash 22.43 15.33 26.24 18.66 22.43 15.33
@@
Nanonets-OCR-s 75.78 62.12 92.07 87.28 75.78 62.12
Gemma-3 27B 75.43 48.19 92.01 75.53 76.16 49.41
LLaMA-4 (Maverick) 17B 47.25 24.73 84.41 68.83 47.25 24.73
GPT-4o 82.43 63.76 94.55 89.14 82.00 64.14
Chitrapathak 34.56 16.10 82.33 68.51 34.56 16.10
Gemini-2.5 Pro 48.48 39.10 67.50 58.76 56.36 47.33
Gemini-2.5 Flash 26.33 11.93 27.17 12.80 27.16 12.84
Table 4: ANLS (Average Normalized Levenshtein Similarity) across English and 11 Indic languages for different models. We denote word- and character-level scores, respectively for each of the 3 separate conditions (lower better). Other variants like WER/CER metrics in Table 7 in Appendix. Gemini-2.5 Flash performs best for Indic languages.
Evaluation on free-form OCR: Table 4 reports Average Normalized Levenshtein Similarity (ANLS) at both word and character levels (lower better) for ORB-En-SynthDog and multilingual ORB-Indic datasets under the three conditions. On free form OCR, docTR and Gemma-3 suffer significant performance drops under rotation, with ANLS rising from 56.2 to 88.0 and 61.6 to 75.1, respectively. In contrast, Gemini-2.5 Flash remains highly robust, increasing only from 22.4 to 26.2, outperforming all other models in the rotated setting, while Gemini-2.5 Pro shows a modest 3-point improvement. Nanonets performs exceptionally best on upright and corrected images but degrades sharply under rotation (2.6 to 44.1), indicating high sensitivity to misalignment and potential training data contamination (roberts2023data; deng2024unveiling). TrOCR and Tesseract are omitted from this comparison where the former is handwriting-specific, and the latter is rule-based, both yielding poor results on this dataset.
On ORB-Indic, Gemini-2.5 Flash consistently delivers the strongest results across all settings, achieving a word-level ANLS of 26.3 on upright documents and showing only a slight increase to 27.2 under rotation. In contrast, Gemini-2.5 Pro performs notably worse than Flash, though it exhibits an 11-point improvement when rotation correction is applied. Additionally, we also include a closed-source OCR solution Chitrapathak [^3] https://bit.ly/chitrapathak which is tailored specifically for Indic languages. Overall, all the models in different categories show poor performance in multilingual settings, with values deteriorating further when documents are rotated. In addition to ANLS, we report WER/CER metrics in Table 7 in the Appendix, following a similar trend. We also show a percentage similarity score in Table 8 in Appendix with both character and word level similarity for completeness. Gemini-2.5 Flash is the most rotation-robust model across both the Indic datasets, while Chitrapathak ranks second, outperforming all the other baselines including LLaMA-4 and Gemma-3.
Failure Analysis: We manually inspect failure cases of our rotation module across both ORB-En and ORB-Indic benchmark datasets. In the ORB-Indic test set, 24 images mostly in Tamil, Hindi, and Telugu were misclassified, primarily confusing $90^{\circ}$ and $0^{\circ}$ with $90^{\circ}$ or $180^{\circ}$, respectively. These errors were typically due to the text being not a majority of the document as well as non uniform placement of the text in the document. In contrast, only 2 misclassifications occurred in the ORB-En-SROIE, where $90^{\circ}$ samples were predicted as $90^{\circ}$. This is primarily due to excessive padding around the text regions as well as off-centered text. However, we observed 25 similar misclassifications in the ORB-En-SynthDog subset, indicating a higher sensitivity to synthetic document variability as well as skewed placement of the text in the document. Figure 3 presents visual examples of these failures, highlighting model confusion caused by unbalanced margins, uneven or extended padding around text regions, and non-standard document layouts.
@@
Figure 3: We show failure cases from the ORB-Indic (a), ORB-En-SynthDog (b) and ORB-En-SROIE datasets (c), discussed in Section 6.
@@ 7 Conclusion
This work highlights the critical importance of rotation correction in OCR pipelines, demonstrating up to 4x improvement in downstream performance when integrated with traditional OCR systems. To rigorously assess model robustness, we introduce OCR-Rotation-Bench (ORB) with 1157 images, comprising (i) ORB-En, derived from rotation-transformed English datasets (both structured and free form), and (ii) ORB-Indic, a multilingual set of 328 real-world documents across 11 Indic languages with ground truth annotations. We present a lightweight rotation classifier built on the Phi-3.5 vision encoder, achieving near-perfect accuracy. Looking ahead, future extensions include arbitrary-angle rotation, skew correction, and layout-aware OCR for handwritten and irregular documents.
@@
Our evaluations reveal significant limitations in SOTA VLMs on basic vision tasks like rotation classification, despite their broader capabilities. Ultimately, our findings underscore the need for future progress to be anchored in well-defined tasks, realistic datasets, and rigorous evaluation protocols for Document Intelligence that accurately reflect real-world challenges.
8 Limitations
Our study presents a focused exploration of rotation correction within OCR pipelines and introduces the OCR-Rotation-Bench (ORB) benchmark. However, some limitations and open directions remain.
Scope and generalization. The current model is trained separately for English and Indic scripts, as cross-script generalization remains challenging. The English classifier, trained on Latin-script documents, achieved only 77.4% accuracy on ORB-Indic, highlighting the influence of script-specific features. Conversely, the Indic model generalizes well across 11 distinct scripts, demonstrating the promise of multilingual training. Future work will pursue a unified, language-agnostic rotation classifier by leveraging visual layout cues such as alignment, borders, and whitespace. We also plan to conduct a human evaluation in future work to quantify reductions in hallucinations and repetitions introduced by rotation errors.
Data and task coverage. ORB primarily includes printed documents and considers four practical discrete rotations ($90^{\circ}$, $0^{\circ}$, $90^{\circ}$, and $180^{\circ}$), without addressing arbitrary angles or skew. We rotated the images by sampling orientations uniformly across the defined rotation classes. However, this study is limited to English and Indic languages; extending to other script families (e.g., European, East Asian or Arabic) is an important future step. Data was collected from open, text-centric document sources (e.g., Wikisource) that might encode content or formatting biases from the originals. We also plan to study diversity of document images in terms of image quality, layout variations, content, and geometric distortions in the future iterations.
Ethical and human aspects. No personally identifiable information (PII) or sensitive content was used. Dataset creation did not involve human participants beyond internal quality verification, and thus did not require IRB approval. All reused datasets and tools are properly cited and used under their respective licenses.
Reproducibility and artifact release. We commit to releasing all benchmark artifacts, models, and evaluation scripts with detailed documentation. Hyperparameters, computing infrastructure, and evaluation protocols are recorded to ensure replicability. We will release the dataset under a license compatible with the original licenses of the datasets used in our work, including SROIE [^4] https://rrc.cvc.uab.es/?ch=13 (CC BY 2.0), SynthDog [^5] https://github.com/clovaai/donut/blob/master/LICENSE (MIT) and Wikisource [^6] https://en.wikisource.org/wiki/Wikisource:Copyright_policy (Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) and the GNU Free Documentation License (GFDL)). Appropriate attribution and share-alike requirements will be respected.
Computational and experimental constraints. We used publicly available pretrained models (e.g., Phi-3.5 Vision Encoder) with moderate GPU resources; large-scale pretraining or fine-grained hyperparameter sweeps were not explored due to compute limits. Inference efficiency can be further improved through optimized deployment pipelines. We compute macro averages in reporting the metrics based on a single model run. While the gains for closed-source models remain modest in some cases, our findings highlight the importance for open-source systems, where this modular architecture offers an impact for advancing robust document understanding. We ensured that the training/validation set is completely disjoint from the test set to ensure there is no data contamination. The training set comprises of in-house data, completely separate from the data used in the benchmark. In the future, we plan to include other SOTA vision encoders as part of our model development.
Potential risks and AI assistance disclosure. We do not anticipate significant ethical risks, but open release of OCR datasets may be misused for large-scale scaling of text extraction from copyrighted or sensitive documents. We will mitigate this to some extent by including clear licensing terms. AI tools (e.g., ChatGPT) were used to refine text writing and section phrasing. All substantive experimental and analytical contributions are human-authored.
Acknowledgements
The authors sincerely thank the leadership at Krutrim and Ola Electric for their support in carrying out this research. We would also like to thank the AI Research team at Krutrim for their valuable suggestions and insightful discussions during different stages of this project.
Appendix
Appendix A Implementation Details
A.1 Model Architecture

### v2

@@ 5 Experimental Design
VLM Rotation Classifier prompt:
Classify this image into one of 12 rotation classes. If the image is not rotated and upright, classify it as 0. If the image is rotated clockwise by 30 degrees, classify it as 30. If it is rotated clockwise by 60 degrees, classify it as 60. If it is rotated clockwise by 90 degrees, classify it as 90. If it is rotated clockwise by 120 degrees, classify it as 120. Continue similarly for 150, 180, 210, 240, 270, 300, and 330 degrees. Reply with just the class number (0, 30, 60, 90, 120, 150, 180, 210, 240, 270, 300, or 330) and nothing else.
English and Multilingual OCR prompt:
You are an OCR engine. You job is to extract the exact text from the given image. Preserve the text formatting, including line breaks, spaces, and any symbols or special characters. Respond in plain text format.
In this section, we detail the experimental strategy used to evaluate both the intrinsic accuracy and downstream impact of the proposed rotation correction module. More technical details are provided in the Appendix 10.
Baselines: We benchmark against both conventional OCR models and Vision-Language Models (VLMs) including docTR [34], transformer-based TrOCR [33] and more recent VLMs including open-weights Gemma-3 (27B) [12], Nanonets-OCR-s [41], LLaMA-4-17B (Maverick)  [54] supporting Indian languages, along with closed-source GPT-4o [24] and Gemini-2.5 Pro and Flash [51]. All prompts are represented in Table 2.
Evaluation metrics: We report field-level accuracy on SROIE and free-form OCR on SynthDog under four rotation angles (0°, 90°, 180°, 270°). For ORB-Indic and the English FUNSD dataset, we report results under a 12-class rotation evaluation. We adopt the evaluation strategy used in prior and contemporary studies [15, 4], reporting Average Normalized Levenshtein Similarity (ANLS) as an interpretable metric  [9] in addition to other variants like Word Error Rate (WER) and Character Error Rate (CER) [49, 42] based on Levenshtein Distance [32].
Model Ideal $\uparrow$ w/o Rotation $\uparrow$ w/ Rotation $\uparrow$
@@
Table 3: Field-level OCR accuracy (in %) on ORB-En-SROIE with and w/o our rotation pipeline. We observe substantial gains for classical and open-source OCR engines, while large VLMs exhibit mixed robustness to rotation.
  • Ideal Upright Conditions: OCR on upright images with an ideal canonical setup.
@@ 6 Results and Discussion
Model ORB-En ORB-Indic
TrOCR 78.18 76.29
docTR 10.00 17.39
Nanonets-OCR-s 8.00 21.53
Gemma-3 27B 16.00 18.74
LLaMA-4 (Maverick) 17B 32.00 29.19
GPT-4o 34.00 17.18
Gemini-2.5 Flash 22.00 23.63
Gemini-2.5 Pro 24.00 22.05
Ours 98.00 96.71
Table 4: Accuracy (%) for 12-class document rotation classification on ORB-En and ORB-Indic. Best in bold, second-best underlined. Conventional models like TrOCR and docTR were finetuned for rotation classification while others are used in 0-shot setting.
Rotation Classification:
Table 4 and Table 7 (Section 11 in Appendix) report the rotation classification accuracy of all evaluated models on our dataset on 2 different cases: 4 class canonical angles (0, 90, 180 and 270 degree rotations) and 12 class rotation problem (0, 30, 60, 90, 120, 150, 180, 210, 240, 270, 300, 330 degree angles). Our proposed classifier, comprising approximately 304M parameters, with 0.5s latency on NVIDIA H100, achieves 96.81% accuracy on ORB-En and 92.68% on the multilingual ORB-Indic dataset on the 4-class classification problem and 98% on ORB-En and 96.71% on ORB-Indic on the 12-class classification problem respectively, demonstrating strong generalization across diverse scripts and layouts. We also conducted 2 ablation studies where dynamic cropping with multi-layer classifier showed the best performance balancing both accuracy and latency shown in Table 6 and Table 8 (Section 11 in Appendix).
Conventional OCR-based systems such as docTR and TrOCR achieve the second-best performance, particularly on the more challenging 12-class rotation task, after being fine-tuned for rotation classification. In contrast, all evaluated VLMs perform poorly, corroborating recent evidence that VLMs struggle on standard computer vision benchmarks [64, 11, 46]. Models such as Gemma-3, LLaMA-4, GPT-4o, and Gemini-2.5 fail to reliably infer document orientation, with GPT-4o marginally outperforming Gemini-2.5 Flash among closed-source models, yet remaining far below our approach. A classical, training-free baseline based on HOG and SIFT orientation statistics yields near-zero exact-match accuracy under the 12-class (30°) setting for both English and Indic datasets, underscoring the limitations of handcrafted methods at fine-grained rotation resolution.
Evaluation on structured OCR: We evaluate the effect of our rotation correction module on downstream OCR using ORB-En-SROIE (Table 3) under the 4-class setting. Field-level accuracy is computed over four equally weighted fields (25% each). Rotation correction yields substantial gains for classical and open-source OCR systems: Tesseract nearly doubles its accuracy, while docTR improves by almost fourfold, underscoring their strong sensitivity to orientation errors. Nanonets-OCR-s attains strong absolute performance and benefits modestly from correction, indicating limited inherent rotation robustness.
Vision–language models show mixed trends. Qwen2.5-VL achieves the best overall accuracy under ideal and corrected settings, with only marginal gains from explicit correction, suggesting partial rotation robustness. In contrast, mPLUG-DocOwl2 performs poorly across all settings. Among closed-source models, Gemini-2.5 Pro benefits noticeably from correction, whereas Gemini-2.5 Flash shows minimal improvement, likely due to internal rotation handling, and GPT-4o remains inferior to conventional OCR engines on structured extraction.
All results are reported without task-specific fine-tuning, highlighting the effectiveness of a lightweight standalone rotation correction module across diverse OCR paradigms.
@@
ORB-En-FUNSD dataset
docTR 38.37 15.02 93.48 75.97 39.86 16.60
Nanonets-OCR-s 99.38 99.23 99.58 99.23 99.41 99.21
Gemma-3 27B 37.41 25.57 73.81 60.36 38.84 26.97
LLaMA-4 (Maverick) 17B 54.44 32.43 83.46 67.18 54.98 33.82
Qwen2.5-VL 31.66 16.72 63.68 51.28 33.08 18.20
GPT-4o 38.05 37.98 79.91 72.68 38.05 37.98
Gemini-2.5 Pro 46.49 35.82 62.81 51.44 47.23 36.53
Gemini-2.5 Flash 34.98 11.93 53.68 39.33 33.26 19.42
@@
Nanonets-OCR-s 91.68 86.92 97.21 95.58 91.68 86.92
Gemma-3 27B 70.38 45.47 95.67 80.50 71.61 47.29
LLaMA-4 (Maverick) 17B 46.68 23.88 92.99 79.50 47.01 24.56
GPT-4o 83.36 65.51 97.70 93.53 83.25 65.79
Chitrapathak 34.05 15.81 93.53 80.40 35.43 17.57
Gemini-2.5 Pro 41.97 27.26 81.49 71.67 51.51 37.83
Gemini-2.5 Flash 35.02 16.30 66.36 49.53 53.66 36.27
Table 5: ANLS (Average Normalized Levenshtein Similarity) under 12-class rotation across English and Indic datasets for different models. Word- and character-level scores are reported for ideal upright inputs, OCR without rotation correction, and OCR with rotation correction (lower is better).
Evaluation on free-form OCR: Table 5 and Table 9 report word- and character-level Average Normalized Levenshtein Similarity (ANLS; lower is better) across English and Indic benchmarks under three settings: ideal upright documents, rotated documents without correction, and rotated documents with rotation correction, for both 4-class and 12-class rotation problems.
On ORB-En-SynthDog (4-class), most OCR systems and VLMs exhibit strong orientation sensitivity. docTR and Gemma-3 degrade sharply under rotation, with word-level ANLS increasing from 56.2 to 88.0 and from 61.6 to 75.2, respectively. Nanonets-OCR-s, despite near-optimal upright performance, degrades severely under rotation (2.6 to 44.1), indicating pronounced orientation bias [47, 13]. VLM robustness is mixed: Gemini-2.5 Flash remains largely invariant (22.4 to 26.2) and performs best without correction, Gemini-2.5 Pro benefits moderately from correction, Qwen-2.5-VL shows partial gains, and mPLUG-DocOwl2 performs poorly across all settings. TrOCR and Tesseract are omitted as unsuitable for free-form documents.
The challenge is further amplified on ORB-En-FUNSD (12-class), which introduces finer-grained orientation variability on unseen documents. All models degrade without correction; however, Gemini-2.5 Flash attains the lowest ANLS under rotation correction, reducing word-level error by nearly 20 points (53.68→33.26). Qwen-2.5-VL shows competitive gains but remains inferior to Gemini-2.5 Flash. docTR, Gemma-3, and LLaMA-4 exhibit compounded errors, underscoring the difficulty of fine-grained orientation handling for end-to-end OCR systems.
On the multilingual ORB-Indic benchmark (966 images across 11 Indic languages), fine-grained orientation effects are amplified in the 12-class setting. Under upright conditions, Chitrapathak achieves the lowest ANLS (34.05 / 15.81), followed closely by Gemini-2.5 Flash (35.02 / 16.30), establishing them as the strongest performers. In contrast, general-purpose OCR and VLM-based systems incur substantially higher errors. Notably, Gemini-2.5 Flash remains the most robust under uncorrected rotation (66.36 / 49.53), despite an almost twofold increase in word-level ANLS. Overall, these results expose a clear gap between Indic-specialized and general-purpose models under fine-grained rotation and highlight the difficulty of multilingual OCR without explicit orientation handling.
Critically, explicit rotation correction is far more effective in the 12-class setting. While several models exhibit partial invariance to coarse (4-class) rotations, fine-grained orientation variability causes consistent degradation across all OCR pipelines when correction is omitted. Incorporating our lightweight rotation module yields substantially larger gains than in the 4-class case, particularly on ORB-En-FUNSD and ORB-Indic (12-class), where models recover a significant portion of upright performance.
Overall, these results indicate that implicit orientation robustness in large OCR and vision–language models is insufficient under fine-grained rotations. Decoupling orientation estimation from recognition consistently improves downstream OCR accuracy, especially for multilingual and free-form documents where visual and linguistic complexity amplifies orientation sensitivity. Additional WER and CER results are reported in Table 10 (Appendix), with percentage-based similarity scores in Table 11. Across all settings, Gemini-2.5 Flash remains the most rotation-robust OCR model on both English and Indic benchmarks.
Ablation study: We also performed an ablation study analyzing three variants of our model in Table 6:
Our ablation study highlights the individual contributions of dynamic cropping and classifier complexity to overall model performance. The baseline variant with a single-layer classifier and dynamic cropping achieves moderate accuracy, indicating that while spatial diversity in input patches is beneficial, a shallow classification head may under utilize the rich representations from the Phi-3.5 encoder. Replacing it with a deeper multi-layer classifier improves results across both ORB-English and ORB-Indic datasets, even without cropping, suggesting that classifier depth significantly enhances discriminative capacity.
Finally, combining both techniques i.e., dynamic cropping and a multi-layer classifier achieves the highest accuracy on both benchmarks (96.81% on ORB-En and 92.68% on ORB-Indic) on the 4-class problem, demonstrating that dynamic cropping provides complementary spatial context that further boosts performance when paired with a sufficiently expressive classifier. This underscores the importance of architectural synergy in maximizing the rotation classification capability of vision-language models.
Model Variant ORB-En (%) ORB-Indic (%)
Single-layer classifier (w/ dynamic cropping) 90.66 82.01
Multi-layer classifier (w/o dynamic cropping) 92.65 86.59
Multi-layer classifier with dynamic cropping (Ours) 96.81 92.68
Table 6: Ablation results comparing classifier complexity and dynamic cropping on ORB-English and ORB-Indic for 4 class classification problem.
@@
Figure 3: We show failure cases from the ORB-En-SynthDog (a), ORB-En-FUNSD (b) and ORB-Indic datasets (c), discussed in Section 6.
Failure Analysis: We manually inspect failure cases of our rotation module across both ORB-En and ORB-Indic benchmark datasets. In the ORB-Indic test set, 45 images, mostly in Tamil, Hindi, and Telugu were misclassified, primarily confusing $90^{\circ}$, $0^{\circ}$ and $240^{\circ}$ with $270^{\circ}$, $180^{\circ}$ or $120^{\circ}$, respectively. These errors were typically due to the text being not a majority of the document as well as non uniform placement of the text in the document.
In contrast, only 2 misclassifications occurred in the ORB-En-SROIE and 1 in ORB-En-FUNSD, where $90^{\circ}$ and $240^{\circ}$ samples were predicted as $270^{\circ}$ and $0^{\circ}$. This is primarily due to excessive padding around the text regions as well as off-centered text. However, we observed 25 similar misclassifications in the ORB-En-SynthDog subset, indicating a higher sensitivity to synthetic document variability as well as skewed placement of the text in the document. Figure 3 presents visual examples of these non-standard document layouts, highlighting model confusion caused by unbalanced margins, uneven or extended padding around text regions.
@@ 7 Conclusion
This work highlights the importance of rotation correction in OCR pipelines, demonstrating up to 4× improvement in traditional OCR systems. To rigorously assess models, we introduce OCR-Rotation-Bench (ORB) with 1863 images, comprising (i) ORB-En, derived from rotation-transformed English datasets (structured and free-form), and (ii) ORB-Indic, a multilingual set of real-world documents across 11 Indic languages with ground-truth annotations. We present a lightweight rotation classifier built on the Phi-3.5 vision encoder, achieving near-perfect accuracy. Future extensions include arbitrary-angle rotation, additional Indic documents, a unified multilingual model, and layout-aware OCR for handwritten and irregular documents.
@@
Our evaluations reveal significant limitations in SOTA VLMs on basic vision tasks like rotation classification, despite their broader capabilities. These results suggest that specialized lightweight models can remain highly competitive for fundamental document preprocessing tasks. Furthermore, our findings highlight the importance of benchmarking simple yet critical tasks that are often overlooked in large-scale multimodal model evaluations. Ultimately, our findings underscore the need for future progress grounded in well-defined tasks, realistic datasets, and rigorous evaluation protocols for Document Intelligence. We hope ORB and our benchmarking protocol will serve as a foundation for future research on robust document orientation and preprocessing in multilingual OCR pipelines.
Supplementary Material
8 Ethics Statement
This work centers on the responsible creation of a benchmark for evaluating OCR robustness to rotation across English and Indic languages. All data used in this study originates from publicly available sources such as Wikisource and established OCR datasets (e.g., SROIE, SynthDog, FUNSD). We ensured that no sensitive or personally identifiable information (PII) was collected or included. The Indic subset, derived from Wikisource, contains only publicly released literary and historical documents that are already digitized and verified. As the research involves only publicly available and non-personal data, formal IRB approval was not required. All external datasets, libraries, and tools are appropriately credited through citations. We exercised care to ensure fairness across scripts and languages, though residual bias may persist due to unequal digital representation among Indic languages. The resulting OCR-Rotation-Bench (ORB) aims to support the development of robust, multilingual, and rotation-aware OCR systems for broader real-world document understanding.
9 Reproducibility Statement
10 Implementation Details
10.1 Model Architecture